A compound containing only C, H, and O, was extracted from the bark of the sassafras tree. The combustion of 63.3 mg produced 172 mg of CO2 and 35.2 mg of H2O. The molar mass of the compound was 162 g/mol. Determine its empirical and molecular formulas.
1
Expert's answer
2020-06-04T10:29:10-0400

12 x172/44 = 47 mg C.

2 x 35.2/18 = 4 mg H.

63.3 - 47-4 = 12.3 mg O.

47/12: 4/1: 12.3/16 = 4 : 4 : 0.76875./0.76875 = 5:5:1.

C5H5O.

12 x 5 + 1 x 5 + 16 x 1 = 81.

162/81=2.

C10H10O2.